Find the general solution of each differential equation, expressing y explicitly in terms of x.

5) dy/dx = x(y+1)


1
Expert's answer
2021-09-15T02:58:35-0400
"dy\/dx=x(y+1)"

"\dfrac{dy}{y+1}=xdx"

Integrate

"\int \dfrac{dy}{y+1}=\int xdx"

"\ln |y+1|=\dfrac{x^2}{2}+\ln C"

"y=Ce^{(x^2\/2)}-1"